# Reduced Order Hysteretic Magnetization Model for Composite Superconductors

### v1

@@ 4.1 S cell
If the irreversibility parameter $\kappa$ is constant, the S cell can be solved with an explicit update rule, as proposed in [27]. Let $\boldsymbol{h}_{\text{rev(p)}}$ be the reversible field computed at the previous time step. For a magnetic field $\boldsymbol{h}$, the updated reversible field $\boldsymbol{h}_{\text{rev}}$ reads:
@@
If the irreversibility parameter $\kappa$ is not constant but a function of $\boldsymbol{b}$ as proposed in Section 3.4 to account for field-dependent $j_{\text{c}}$, the update rule is no longer explicit. In such a case however, we observed that a good approximate solution is found easily in a few fixed point iterations, that is, by solving successively Eq. (4.1) with updated values of $\kappa(\boldsymbol{b})=\kappa(\mu_{0}\boldsymbol{h}_{\text{rev}})$ until $\boldsymbol{h}_{\text{rev}}$ does no longer change significantly.
It is remarkable that the implementation of the S cell entirely consists in the simple test of Eq. (4.1). It only requires the knowledge of the previous reversible field, which has to be stored as an internal variable in the numerical implementation.
@@ 4.2 CS cell
For the CS cell described in Section 3.3, the approach is similar. However, two tests are now necessary to account for the two dry friction elements. We define
@@
and denote by $\boldsymbol{g}_{\text{(p)}}$ its value at the previous time step. We start by updating $\boldsymbol{g}$, as a function of the new $\boldsymbol{h}$ as follows:
@@
with $\boldsymbol{h}_{\text{rev(p)}}$ the reversible field at the previous time step and $\Delta t$ the current time step. Solving this equation for $\boldsymbol{h}_{\text{rev}}^{\text{trial}}$, and then evaluating $\boldsymbol{h}_{\text{coupling}}^{\text{trial}}=\tau_{\text{c}}\dot{\boldsymbol{h}}_{\text{rev}}^{\text{trial}}$ gives the trial value (we drop the $\approx$ sign)
@@
All instantaneous power quantities can be computed by Eqs. (12), (13), (17), (20), and (21).
@@
The simulation of the CS cell requires to save the values of vectors $\boldsymbol{g}$ and $\boldsymbol{h}_{\text{rev}}$ at each time step.
@@ 4.3 Chain of cells
In a hysteresis model with a chain of cells as proposed in Section 3.5, each cell $k$ is subject to the same magnetic field $\boldsymbol{h}$ and produces a distinct magnetic flux density $\boldsymbol{b}_{k}$. The total magnetic flux density $\boldsymbol{b}$ is computed as a weighted sum of the $\boldsymbol{b}_{k}$ as defined in Eq. (23). This is the only additional step compared to the single cell case.
@@ 4.4 Inclusion in a finite element $\phi$-formulation
Let us consider a numerical domain $\Omega$. It is decomposed into the superconducting strand (or any other superconducting system) to homogenize, denoted as $\Omega_{\text{m}}$, and the complementary domain, denoted as $\Omega_{\text{m}}^{\text{C}}$. For simplicity, we assume that $\Omega_{\text{m}}^{\text{C}}$ is non-conducting, but the approach can be easily extended to conducting domains as well. Also, in this paper, we only consider cases with no transport current in $\Omega_{\text{m}}$.
@@
These equations can be solved numerically with the finite element method, on a mesh, i.e., a spatial discretization of the numerical domain $\Omega$.
As the hysteresis law is driven by the magnetic field $\boldsymbol{h}$, it is more natural to consider an $h$-conform formulation, such as the $h$-$\phi$-formulation [41, 42]. In this particular case with no net currents in $\Omega_{\text{m}}$ and a non-conducting $\Omega_{\text{m}}^{\text{C}}$, this formulation can be reduced to a $\phi$-formulation [29].
Starting from an initial solution, it consists of finding a magnetic field $\boldsymbol{h}=\textrm{{grad}}\,\phi$ with $\phi\in\Phi(\Omega)$ such that, at subsequent time instants, $\forall\boldsymbol{h}^{\prime}=\textrm{{grad}}\,\phi^{\prime}$ with $\phi^{\prime}\in\Phi_{0}(\Omega)$,
@@
with $\Phi(\Omega)$ and $\Phi_{0}(\Omega)$ appropriate function spaces, and where the notation $\left(\boldsymbol{v}\,,\boldsymbol{w}\right)_{\Omega}$ denotes the integral over $\Omega$ of the dot product of any two vector fields $\boldsymbol{v}$ and $\boldsymbol{w}$.
The presence of the hysteresis law in Eq. (37) makes the system history-dependent. The fields $\boldsymbol{h}_{\text{rev}}$ and $\boldsymbol{g}$ must therefore be saved at each steps. They are defined in $\Omega_{\text{m}}$ only and are chosen to be element-wise constant vectors.
The hysteresis law also makes the system nonlinear. An iterative scheme is therefore necessary for a numerical simulation. In this case, we observed that the Newton-Raphson method leads to efficient simulations. From an initial iterate $\boldsymbol{h}_{0}$, it consists in solving successively a linearized version of Eq. (37) until a given convergence criterion is met. At iteration $i\in\mathbb{N}_{0}$, the formulation reads, in terms of the unknown field $\boldsymbol{h}_{i}$,
@@ 5 Parameter Identification
The hysteresis model presented in Section 3 provides a flexible tool able to capture a variety of different hysteretic responses thanks to the approach with several cells connected as a chain. The parameter values of each of these cells must be properly chosen in order to reproduce reference solutions or measurement data.
@@ 5.1 Chain of S cells
In a chain of S cells, there are $2N$ parameters to be defined: the constant weights $\alpha_{k}$ and the irreversibility parameters $\kappa_{k}(\boldsymbol{b})=f_{\kappa}(\boldsymbol{b})\bar{\kappa}_{k}$. The number of cells $N$ must also be chosen as a trade-off between computational cost, implementation easiness and accuracy.
As a reference solution, we consider the magnetization curve of a superconducting strand subject to a unidirectional field $\boldsymbol{h}_{\text{app}}(t)=h_{\text{app}}(t)\hat{\boldsymbol{e}}_{y}=h_{\text{max}}\sin(2\pi ft)\hat{\boldsymbol{e}}_{y}$, with a sufficiently large field $h_{\text{max}}$ for the system to be fully penetrated, and at a sufficiently low frequency $f$ for the coupling and eddy current effects to be negligible. We compute the associated average magnetic flux density in the strand $\boldsymbol{b}_{\text{in}}$ as
@@
using Eqs. (1) and (7). An example of such a curve for the composite strand described in Section 2 is given in Fig. 15, where $\mu_{0}h_{\text{max}}=2$ T. The curve is obtained with a standard 2D $h$-$\phi$-formulation that does not model coupling currents between the filaments, at $f=0.01$ Hz.
@@
Step 1: Fitting the field-dependent scaling
@@
For a chain of S cell, using Eqs. (8) and (23), the total magnetization is given by
@@
with $m_{\text{max}}$ (in A/m) the maximum magnetization. As a result, the scaling function $f_{\kappa}(\boldsymbol{b})$ directly describes the shape of the purple curves in Fig. 15.
A simple approach consists in choosing $f_{\kappa}(\boldsymbol{b})=j_{\text{c}}(\boldsymbol{b})/j_{\text{c}}(\boldsymbol{0})$ as a scaling over the whole field range. This however does not lead to very accurate results at low fields for which $j_{\text{c}}$ strongly depends on $\boldsymbol{b}$. The reason is that the strand magnetization depends on the actual field distribution inside it, which is not uniform, and not only on the average vector $\boldsymbol{b}_{\text{in}}$. The larger the filaments in the strand, the larger the variation of $\boldsymbol{b}$ within each filament.
A better approach is to directly identify the scaling $f_{\kappa}(\boldsymbol{b})$ on the reference curve. The purple curves can therefore be used to directly define $f_{\kappa}(\boldsymbol{b})$ in the available field range. Outside of the field range, $f_{\kappa}(\boldsymbol{b})=j_{\text{c}}(\boldsymbol{b})/j_{\text{c}}(\boldsymbol{0})$ can be used as a first approximation.
@@
Step 2: Choosing the $\bar{\kappa}_{k}$ and fitting the weights $\alpha_{k}$
@@
The proposed procedure consists in choosing a priori the $N$ values $\bar{\kappa}_{k}$ and then fixing the weights $\alpha_{k}$ accordingly. For simplicity, we arrange the $\bar{\kappa}_{k}$ values in increasing order with respect to $k$ and start with $\bar{\kappa}_{1}=0$ A/m.
We consider the virgin magnetization curve represented by the dotted curve in Fig. 15. Initially, $h_{\text{rev},k}=0$, $\forall k=1,\dots,N$. As the magnetic field $h_{\text{in}}$ progressively increases, it reaches the surface of the cell spheres one by one (see Fig. 8). This happens at successive threshold field values that depend on the $\bar{\kappa}_{k}$ constants and on the $f_{\kappa}(b_{\text{in}})$ scaling determined in Step 1. We denote these threshold fields as $h_{k}$.
For $h<h_{k}$, we have $h_{\text{rev},j}=0$, $\forall j=k,\dots,N$, such that only the first $k-1$ cells contribute to the magnetic flux density $b$. The weights $\alpha_{k-1}$ can therefore be successively determined by forcing the solution of the hysteresis model to match the reference solution at each $h_{k}$, for $k=2,\dots,N$. The last weight $\alpha_{N}$ is calculated so that all weights add up to one.
The choice of the number of cells $N$ and the a priori distribution of values for the $\bar{\kappa}_{k}$ depend on the sought accuracy and application. This is discussed in Section 6.
@@ 5.2 Chain of CS cells
For a chain of CS cells, there are $5N$ parameters to choose: $\alpha_{k}$, $\kappa_{k}$, $\tau_{\text{e},k}$, $\tau_{\text{c},k}$, and $\chi_{k}$, for $k=1,\dots,N$. To identify all these parameters, reference solutions at different frequencies are necessary.
A reference solution at a sufficiently low frequency for the eddy and coupling current effects to be negligible can be used exactly as described in the previous section to identify the $\alpha_{k}$ and $\kappa_{k}$ (with both the $f_{\kappa}(\boldsymbol{b})$ scaling and the $\bar{\kappa}_{k}$ values). This leads to a number of $3N$ remaining parameters.
The time constants $\tau_{\text{e},k}$ and $\tau_{\text{c},k}$ can be chosen so that the positions of the peaks in eddy and coupling losses respectively correspond to those of reference solutions, such as shown in Fig. 3. Choosing identical values for all the cells already leads to good results within limited amplitude ranges. To reproduce the observation that the maxima in coupling and eddy loss are observed at lower frequencies when the field amplitude increases, one can choose larger values of $\tau_{\text{e},k}$ and $\tau_{\text{c},k}$ for cells associated with larger irreversibility parameters $\kappa_{k}$. This will be illustrated in Section 7.
Once the time constants are chosen, the only remaining parameters are the irreversibility parameters $\chi_{k}(\boldsymbol{b})=f_{\chi}(\boldsymbol{b})\bar{\chi}_{k}$. The scaling $f_{\chi}(\boldsymbol{b})$ can be chosen as in the static case (Step 1), based on a reference solution at a sufficiently high frequency for the dynamic effects to be visible in the magnetization curve, e.g., at $f=100$ Hz for the 54-filament strand, as shown in Fig. 7. Then, the values $\bar{\chi}_{k}$ can be identified in order to best reproduce the transition branches, as in Step 2 of the static case, but now with values $\bar{\chi}_{k}$ as unknowns instead of the weights $\alpha_{k}$ which are already fixed.
@@ 6 Results with a chain of S cells
(a) Major loop $h_{\text{app}}^{(1)}(t)$.
(b) Minor loops $h_{\text{app}}^{(2)}(t)$.
Figure 16: Comparison of the hysteresis model results with a reference solution. Unidirectional transverse magnetic field. Chain of S cells with $N=6$ cells, parameters of Table 2. (a) Harmonic excitation $h_{\text{app}}^{(1)}(t)$. (b) Biharmonic excitation $h_{\text{app}}^{(2)}(t)$. (Up) Magnetization loop. (Middle) Dissipated power per unit length as a function of time, equal to $a\,p_{\text{irr}}(t)$, with $a=\pi d^{2}/4$ the surface area of the strand cross section. (Down) Internal magnetic field, the input of the hysteresis model.
The first application consists in the 54-filament strand presented in Section 2, subject to a transverse field varying sufficiently slowly for the eddy and coupling current effects to be negligible, that is, with frequencies of the order of $f=0.01$ Hz. For the reference solution, we assume a non-conducting matrix and use a classical $h$-$\phi$-formulation [41], so that coupling effects are completely removed.

### v2

@@ 4.1 S cell
If the irreversibility parameter $\kappa$ is constant, the S cell can be solved with an explicit update rule, as proposed in [27]. Let $\boldsymbol{h}_{\text{rev(p)}}$ be the reversible field computed at the previous time step, as a result of the history of $\boldsymbol{h}$. For a magnetic field $\boldsymbol{h}$ at the new time step, the reversible field $\boldsymbol{h}_{\text{rev}}$ is updated as follows:
@@
If the irreversibility parameter $\kappa$ is not constant but a function of $\boldsymbol{b}$ to account for field-dependent $j_{\text{c}}$, the update rule is no longer explicit. In such a case, an approximate solution can be computed easily with a few fixed point iterations, that is, by solving successively Eq. (4.1) with updated values of $\kappa(\boldsymbol{b})=\kappa(\mu_{0}\boldsymbol{h}_{\text{rev}})$ until $\boldsymbol{h}_{\text{rev}}$ does no longer change significantly.
The implementation of the S cell is very simple, as it entirely consists in the test of Eq. (4.1). It only requires the knowledge of the previous reversible field $\boldsymbol{h}_{\text{rev}}$, which has to be stored as an internal variable during the numerical simulation.
@@ 4.2 CS cell
For the CS cell described in Section 3.3, the approach is similar. However, two tests are now necessary to account for the two dry friction elements. With
@@
and $\boldsymbol{g}_{\text{(p)}}$ its value at the previous time step. We start by updating $\boldsymbol{g}$, as a function of the new $\boldsymbol{h}$ as follows:
@@
with $\boldsymbol{h}_{\text{rev(p)}}$ the reversible field at the previous time step and $\Delta t$ the current time step size. Solving this equation for $\boldsymbol{h}_{\text{rev}}^{\text{trial}}$, and then evaluating $\boldsymbol{h}_{\text{coupling}}^{\text{trial}}=\tau_{\text{c}}\dot{\boldsymbol{h}}_{\text{rev}}^{\text{trial}}$ gives the trial value (we drop the $\approx$ sign)
@@
The instantaneous power quantities can be computed by Eqs. (12), (13), (18), (21), and (22).
@@
The implementation of the CS cell is straightforward. It only consists in two tests and a few update equations. The simulation of the CS cell requires to keep track of the values of vectors $\boldsymbol{g}$ and $\boldsymbol{h}_{\text{rev}}$ at the previous time step.
@@ 4.3 Chain of cells
In a ROHM model with a chain of cells as proposed in Section 3.4, each cell $k$ is subject to the same magnetic field $\boldsymbol{h}$ and produces a distinct magnetic flux density $\boldsymbol{b}_{k}$. The total magnetic flux density $\boldsymbol{b}$ is then computed as a weighted sum of the $\boldsymbol{b}_{k}$, as defined in Eq. (26). This is the only additional step compared to the single cell case.
@@ 4.4 Inclusion in a finite element $\phi$-formulation
Let us consider a numerical domain $\Omega$. It is decomposed into the superconducting strand (or any other superconducting conductor) to be homogenized, denoted as $\Omega_{\text{m}}$, and the complementary domain, denoted as $\Omega_{\text{m}}^{\text{C}}$. For simplicity, we assume that $\Omega_{\text{m}}^{\text{C}}$ is non-conducting, but the approach can be easily extended to conducting domains as well. Also, in this paper, we only consider cases with no transport current in $\Omega_{\text{m}}$.
@@
These equations can be solved numerically with the FE method, on a mesh, i.e., a spatial discretization of the numerical domain $\Omega$.
As the hysteresis law is driven by the magnetic field $\boldsymbol{h}$, it is easier to consider an $h$-conform formulation, such as the $h$-$\phi$-formulation [39, 40]. In this particular case with no net currents in $\Omega_{\text{m}}$ and a non-conducting $\Omega_{\text{m}}^{\text{C}}$, this formulation can be reduced to a $\phi$-formulation [29].
Starting from an initial solution, it consists in finding a magnetic field $\boldsymbol{h}=\textrm{{grad}}\,\phi$ with $\phi\in\Phi(\Omega)$ such that, at subsequent time instants, $\forall\boldsymbol{h}^{\prime}=\textrm{{grad}}\,\phi^{\prime}$ with $\phi^{\prime}\in\Phi_{0}(\Omega)$,
@@
where the notation $\left(\boldsymbol{v}\,,\boldsymbol{w}\right)_{\Omega}$ denotes the integral over $\Omega$ of the dot product of any two vector fields $\boldsymbol{v}$ and $\boldsymbol{w}$. The space $\Phi(\Omega)$ contains functions in $H^{1}(\Omega)$ that satisfy appropriate essential boundary conditions. The space $\Phi_{0}(\Omega)$ is the same space but with homogeneous essential boundary conditions. These function spaces are discretized with lowest-order node functions [41].
The presence of the hysteresis law in Eq. (40) makes the system history-dependent. The fields $\boldsymbol{h}_{\text{rev}}$ and $\boldsymbol{g}$ must therefore be saved at each time step. They are defined in $\Omega_{\text{m}}$ only and are chosen to be element-wise constant vectors, with no continuity constraint across element interfaces.
The hysteresis law also makes the system nonlinear. An iterative scheme is therefore necessary for a numerical simulation. We observed that the Newton-Raphson method leads to very efficient simulations. From an initial iterate $\boldsymbol{h}_{0}$, it consists in solving successively a linearized version of Eq. (40) until a given convergence criterion is met. At iteration $i$, the formulation reads, in terms of the unknown field $\boldsymbol{h}_{i}$, and the previous (known) iterate $\boldsymbol{h}_{i-1}$,
@@ 5 Parameter Identification
The ROHM model presented in Section 3 provides a flexible tool able to capture a variety of different hysteretic responses thanks to the approach with several cells connected as a chain. The parameter values of each of these cells must be properly chosen in order to reproduce reference solutions.
@@ 5.1 Chain of S cells
In a chain of S cells, there are $2N$ parameters to be defined: the irreversibility parameters $\kappa_{k}(\boldsymbol{b})=f_{\kappa}(\boldsymbol{b})\bar{\kappa}_{k}$ and the constant weights $\alpha_{k}$. The number of cells $N$ must be chosen as a trade-off between computational cost, implementation easiness and accuracy.
As a reference solution, we consider the magnetization curve of a superconducting strand subject to a unidirectional field $\boldsymbol{h}_{\text{app}}(t)=h_{\text{app}}(t)\hat{\boldsymbol{e}}_{y}=h_{\text{max}}\sin(2\pi ft)\hat{\boldsymbol{e}}_{y}$, with a sufficiently large field $h_{\text{max}}$ for the system to be fully penetrated, and at a sufficiently low frequency $f$ for the coupling and eddy current effects to be negligible. Based on the reference solution, we compute the average magnetic flux density in the strand $\boldsymbol{b}_{\text{in}}$ as
@@
using Eqs. (1) and (7). An example of the $\boldsymbol{b}_{\text{in}}$ and magnetization curves for the composite strand described in Section 2 is given in Fig. 14, with $\mu_{0}h_{\text{max}}=2$ T and $f=0.01$ Hz. The solution is obtained with a standard 2D FE $h$-$\phi$-formulation without any filament coupling.
@@
Step 1: Fitting the field-dependent scaling.
@@
For a chain of S cells, using Eqs. (8) and (26), the total magnetization is given by
@@
with $m_{\text{max}}$ (A/m) the maximum magnetization. Therefore, the scaling function $f_{\kappa}(\boldsymbol{b})$ directly describes the shape of the purple curves in Fig. 14, scaled by $m_{\text{max}}$.
As a result, the scaling $f_{\kappa}(\boldsymbol{b})$ can be defined as an interpolation function of the purple curves, scaled by $m_{\text{max}}$. Outside of the field range of the reference solution, $f_{\kappa}(\boldsymbol{b})$ can be extrapolated smoothly, e.g., using the field-dependent critical current density as a scaling. The fact that the model reproduces a maximum magnetization equal to $m_{\text{max}}$ will be ensured in step 2.
@@
Note that choosing $f_{\kappa}(\boldsymbol{b})=j_{\text{c}}(\boldsymbol{b})/j_{\text{c}}(\boldsymbol{0})$ as a scaling over the whole field range, instead of defining it based on the purple curve as proposed above, does not lead to very accurate results, especially at low fields. The reason is that the strand magnetization depends on the actual field distribution inside it, which is not uniform (but unknown for the homogenized ROHM model), and not only on the average vector $\boldsymbol{b}_{\text{in}}$.
Step 2: Choosing the $\bar{\kappa}_{k}$ and fitting the weights $\alpha_{k}$.
@@
The proposed procedure consists in choosing a priori the $N$ values $\bar{\kappa}_{k}$ and then fixing the weights $\alpha_{k}$ accordingly. For simplicity, we arrange the $\bar{\kappa}_{k}$ values in increasing order with respect to $k$ and we start with $\bar{\kappa}_{1}=0$ A/m.
We consider the virgin magnetization curve represented by the green curve starting at origin in Fig. 14. Initially, $h_{\text{rev},k}=0$, $\forall k=1,\dots,N$. As the magnetic field $h_{\text{in}}$ increases, it reaches the surface of the cell spheres one by one (see Figs. 7 and 15). This happens at successive threshold field values that depend on the $\bar{\kappa}_{k}$ constants and on the $f_{\kappa}(b_{\text{in}})$ scaling determined in Step 1. We denote these threshold fields as $h_{k}$, as illustrated in Fig. 16.
Figure 15: Illustration of the magnetic field and a chain of S cells with $N=4$ (the first cell is with $\kappa_{1}=0$ A/m), from virgin state to $\|\boldsymbol{h}\|=h_{4}$. Each sphere (circle) represents one cell. The center of the spheres are at positions $\boldsymbol{h}_{\text{rev},k}$, established by the history of $\boldsymbol{h}$, and their field-dependent radius is equal to $\kappa_{k}(\boldsymbol{b})$. The relative scale is respected compared to Fig. 16. Variations of the magnetic field $\boldsymbol{h}$ that stay inside the smallest non-degenerated sphere produce no loss.
[…]
For $h<h_{k}$, we have $h_{\text{rev},j}=0$, $\forall j=k,\dots,N$, such that only the first $k-1$ cells contribute to the magnetic flux density $b$. The weights $\alpha_{k-1}$ can therefore be consecutively adjusted by forcing the solution of the ROHM model to match the reference solution at each $h_{k}$, for $k=2,\dots,N$. This adjustment can be automated with a simple iterative routine. The influence of the weights $\alpha_{k}$ on the ROHM model solution is illustrated in Fig. 16. The last weight $\alpha_{N}$ is calculated so that all weights add up to one.
The choice of the number of cells $N$ and the a priori distribution of values for the $\bar{\kappa}_{k}$ depend on the desired accuracy and application. This is discussed in Section 6.
@@ 5.2 Chain of CS cells
For a chain of CS cells, there are $5N$ parameters to be defined: $\kappa_{k}$, $\alpha_{k}$, $\tau_{\text{e},k}$, $\tau_{\text{c},k}$, and $\chi_{k}$, for $k=1,\dots,N$. To identify all these parameters, reference solutions at different frequencies are necessary. We decompose the identification procedure in three steps.
Step 1: Reproducing the rate-independent regime.
A reference solution at a sufficiently low frequency, so that the eddy and coupling current effects are negligible, can be used exactly as described in the previous section to define the $\kappa_{k}$ (with both the $f_{\kappa}(\boldsymbol{b})$ scaling and the $\bar{\kappa}_{k}$ values) and $\alpha_{k}$. This leads to a number of $3N$ remaining parameters.
Step 2: Fitting the time constants.
The time constants $\tau_{\text{e},k}$ and $\tau_{\text{c},k}$ can be chosen so that the positions of the peaks in eddy and coupling losses correspond to those of reference solutions, such as shown in Fig. 3. Choosing identical values for all the cells already leads to good results within limited amplitude ranges. To reproduce the observation that the maxima in coupling and eddy loss are observed at lower frequencies when the field amplitude increases, one can choose larger values of $\tau_{\text{e},k}$ and $\tau_{\text{c},k}$ for cells associated with larger irreversibility parameters. This will be illustrated in Section 7.
Step 3: Reproducing the coupled filament magnetization.
Once the time constants are fixed, the only remaining parameters are the $\chi_{k}(\boldsymbol{b})=f_{\chi}(\boldsymbol{b})\bar{\chi}_{k}$. The scaling $f_{\chi}(\boldsymbol{b})$ can be defined as in the static case (Step 1), based on a reference solution at a sufficiently high frequency for the dynamic effects to be visible in the magnetization curve, e.g., at $f=100$ Hz for the 54-filament strand, as shown in Fig. 6. Then, the values $\bar{\chi}_{k}$ can be identified in order to best reproduce the transition branches, as in Step 2 of the static case, but now with values $\bar{\chi}_{k}$ as unknowns instead of the weights $\alpha_{k}$ which are already fixed.
@@ 6 Results with a chain of S cells
The first application consists of the 54-filament strand presented in Section 2, subject to a transverse field varying sufficiently slowly for the eddy and coupling current effects to be negligible. We choose a frequency $f=0.01$ Hz. For the reference solution, we assume a non-conducting matrix and use a classical $h$-$\phi$-formulation [39], so that coupling effects are not present.